OpenStudy (anonymous):

Use the compound interest formula A=P(1+R)^t and the given information to solve for R. A=$2600 P=$1800 T=7

OpenStudy (kropot72):

Plugging in the given values we get: \[2600=1800(1+r)^{7}\] Do you want help step by step to find r?

OpenStudy (kropot72):

@MawyKay Are you there?

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

Do you want to go step by step? The first step is to divide both sides of the equation by 1800. This step will isolate the expression with the exponent.

OpenStudy (anonymous):

1.44=(1+r)^7 correct?

OpenStudy (anonymous):

do you know what the next step is?

OpenStudy (kropot72):

Good work! However take the number on the left hand side to four decimal places, giving 1.4444. Next take logs of both sides giving: \[\log_{} 1.4444=7\log_{} (1+r)\ .............(1)\] and now divide both sides of equation (1) by 7.

OpenStudy (anonymous):

what does that look like? I'm confused.

OpenStudy (kropot72):

\[\log_{} 1.4444=7\log_{} (1+r)\ ...........(1)\] and now divide both sides of equation (1) by 7, giving \[\frac{\log_{} 1.4444}{7}=\log_{} (1+r)\ .......(2)\] 0.0228125=log (1 + r) ..............(3) Therefore by the rules of logs \[10^{0.0228125}=(1+r)\] 1.053932 = 1 + r r = 0.053932 or 5.3932%